OpenStudy (the_puppet1):

To solve the equation 6 = 2 - 4x, Alex decided to subtract 2 from both sides of the equation. Which equation is the result of that step? A. 3 = -4x B. 4 = -4x C. 8 = -4x D. 12 = -4x

OpenStudy (the_puppet1):

@welshfella

Directrix (directrix):

6 = 2 - 4x -2 -2 ------------------

Directrix (directrix):

@The_Puppet1 What results with 2 issubtracted from the equation?

Directrix (directrix):

6 - 2 = ? 2 - 2 = ?

OpenStudy (the_puppet1):

Is 0?

OpenStudy (the_puppet1):

6 - 2 = 4 2 - 2 = 0

Directrix (directrix):

6 = 2 - 4x -2 -2 ------------------ 4 = - 4x

Directrix (directrix):

Do you agree?

OpenStudy (the_puppet1):

Yeah

Directrix (directrix):

Which option is that?

Directrix (directrix):

@The_Puppet1 ^^^